A company uses the formula C(x)= 0.02x^2-3.4x + 150 to model the unit cost in dollars for producing x stabilizer bars. For what number of bars is the unit cost at its minimum? What is the unit cost at that level of production?
Answer by Alan3354(69443)

A company uses the formula C(x)= 0.02x^2-3.4x + 150 to model the unit cost in dollars for producing x stabilizer bars. For what number of bars is the unit cost at its minimum? What is the unit cost at that level of production?
------------------
C(x)= 0.02x^2-3.4x + 150
To find the minimum, set the 1st derivative to 0
0.04x - 3.4 = 0
Multiply by 100
4x = 340
x = 85
-------
Sub 85 for x in the equation
C(x)= 0.02x^2-3.4x + 150
C(85) = 0.02*85*85 - 3.4*85 + 150
= $583.50
That's the cost of 85 bars.
For one, $583.50/85 = $6.865